OpenStudy (anonymous):

perpendicular slope 2

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

find slope of the line

OpenStudy (gorv):

\[slope=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (gorv):

\[=\frac{ 1-(-1) }{ 2-0 }=\frac{ 2 }{ 2 }=1\]

OpenStudy (gorv):

for perpendicular m*1=-1 m=-1=slope of perpendicular

OpenStudy (gorv):

@edjoey